OpenStudy (howard-wolowitz):

Am I correct: I got 3.96240625 which is 3.9624

Nnesha (nnesha):

how did you get that answer ? did you apply change of base formula ?? :-)

OpenStudy (howard-wolowitz):

yes.. but i dont understand

Nnesha (nnesha):

bec there are two method to solve exponential equation

Nnesha (nnesha):

okay what's the change of base formula ?

OpenStudy (howard-wolowitz):

ln(4z)=ln(243)

OpenStudy (howard-wolowitz):

zln(4)=ln(243)

OpenStudy (howard-wolowitz):

2zln(2)=ln(243)

Nnesha (nnesha):

okay so you can divide both side by ln 4

Nnesha (nnesha):

hmm 2 ??

OpenStudy (howard-wolowitz):

2zln(2)=1(5ln(3))

OpenStudy (howard-wolowitz):

no 2?

Nnesha (nnesha):

change of base formula \[\huge\rm log_a b = \frac{ \log b }{ \log a }\] that's it

OpenStudy (howard-wolowitz):

z=5ln(3)//////////2ln(2)

Nnesha (nnesha):

how did you get 5 ? z= 5ln(3) ^ ?

OpenStudy (howard-wolowitz):

here: The left-hand side of the equation is equal to the exponent of the logarithm argument because the base of the logarithm equals the base of the argument. zln(4)=ln(243)

Nnesha (nnesha):

 first of all change to log form